OpenStudy (tiffanymak1996):

Def. 1 : A function f of two variables is said to be differentiable at (x0,y0) provided fx(x0,y0) both exist and [insert equation 1 here] Def 2: A function f of three variables is said to be differentiable at (x0,y0,z0) provided fx(x0,y0,z0), fy(x0,y0,z0) and fz(x0,y0,z0) exist and [insert eq 2 here]. Use the above two def (s) to prove a) a constant function of 2 or 3 variables is differentiable everywhere. b) a linear function of 2 or 3 variables is differentiable everywhere.

OpenStudy (tiffanymak1996):

eq1: \[\lim_{(\Delta x,\Delta y) \rightarrow (0,0)} \frac{ \Delta f - f _{x} (x _{0}, y _{0}) \Delta x - f _{y}(x _{0}, y _{0}) \Delta y}{ \sqrt(\Delta x)^2 + (\Delta y)^2 } =0\]

OpenStudy (tiffanymak1996):

the second eq is very similar, except where it was xy, it is xyz.

OpenStudy (tiffanymak1996):

I think I know how to prove this, I'm just not quite sure what it means by a constant function of 2 or 3 variables.

zepdrix (zepdrix):

Constant function? D: Hmm do they mean like,\[\large f(x,y)=c\]

OpenStudy (tiffanymak1996):

sorry, my computer crashed. :(

OpenStudy (tiffanymak1996):

maybe, thanks, I'll give it a go.

OpenStudy (tiffanymak1996):

yup, definitely. Thanks again.